OpenStudy (anonymous):

Find -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3. -67 -17 -31 1

OpenStudy (anonymous):

sub in the values of a,b and c into the equation -(3)^2 - 3(2)^3 + (-3)^2 +2(2)^3 (simplify) = -(9) - 3(8) + (9) + 2(8) = -9 - 24 + 9 + 16 = -8

OpenStudy (anonymous):

if this answer is wrong, you didn't put correct brackets in your equation so i can't tell if the exponents apply to the number or the negative signs too